# Towards Understanding Valuable Preference Data for Large Language Model Alignment

### v1

@@ Abstract
Large language model (LLM) alignment is typically achieved through learning from human preference comparisons, making the quality of preference data critical to its success. Existing studies often pre-process raw training datasets to identify valuable preference pairs using external reward models or off-the-shelf LLMs, achieving improved overall performance but rarely examining whether individual, selected data point is genuinely beneficial. We assess data quality through individual influence on validation data using our newly proposed truncated influence function (TIF), which mitigates the over-scoring present in traditional measures and reveals that preference data quality is inherently a property of the model. In other words, a data pair that benefits one model may harm another. This leaves the need to improve the preference data selection approaches to be adapting to specific models. To this end, we introduce two candidate scoring functions (SFs) that are computationally simpler than TIF and positively correlated with it. They are also model dependent and can serve as potential indicators of individual data quality for preference data selection. Furthermore, we observe that these SFs inherently exhibit errors when compared to TIF. To this end, we combine them to offset their diverse error sources, resulting in a simple yet effective data selection rule that enables the models to achieve a more precise selection of valuable preference data. We conduct experiments across diverse alignment benchmarks and various LLM families, with results demonstrating that better alignment performance can be achieved using less data, showing the generality of our findings and new methods.
@@ 1 Introduction
Reinforcement learning with human feedback (RLHF) has emerged as a dominant fine-tuning paradigm for aligning large language models (LLMs) with human preferences (Bai et al., 2022b; Wang et al., 2024). Whether through training explicit reward models (Lambert et al., 2024) or optimizing policy with implicit rewards (Rafailov et al., 2023), the success of RLHF hinges heavily on the availability of high-quality preference data. Previous works (Shen et al., 2024; Pattnaik et al., 2024; Morimura et al., 2024; Deng et al., 2025; Chen et al., 2024) typically leverage external reward models or off-the-shelf LLMs to filter raw data, treating selected data as reliable sources for training their models or releasing them for open-source use (Cui et al., 2023; Bai et al., 2022a). This pre-processing paradigm dominates the RLHF community to improve data quality, contributing remarkably to the success of many publicly available LLMs such as Llama (Vavekanand & Sam, 2024), Qwen (Yang et al., 2025) and DeepSeek (Guo et al., 2025).
However, such pre-processing implicitly assumes that quality is an intrinsic property of data themselves: regardless of training configurations or models, certain data are consistently presumed to be more valuable for alignment than others. Therefore, a seemingly more realistic perspective is that the data quality is also a property of the model (Grosse et al., 2023; Xia et al., 2024). That is to say, some data points may be beneficial for certain models or configurations while being detrimental at others. We validate this new assumption as more reasonable based on the influence function (IF) (Koh & Liang, 2017), which quantifies the impact of each training data point on validation performance, thereby reflecting data quality. Nevertheless, we observe that the influence scores may overfit to the validation data, an issue that is particularly severe for open-world models like LLMs. To address this, we propose and verify a simple modification to the original IF, namely truncated influence function (TIF), which is shown to be more reliable for preference data selection and results in better overall performance. Using TIF, we verify that data quality is model-dependent, varying across different models, with some data points proving beneficial for certain models yet harmful at others.
@@ 2 Preliminary
Direct Preference Alignment. Traditional RLHF (Bai et al., 2022b) often follows a two-stage pipeline: first training a reward model on preference data, and then optimizing the policy LLM using reinforcement learning (RL) algorithm such as PPO (Schulman et al., 2017). The two-stage RLHF is relatively complicated and resource-intensive. Recently, many studies (Rafailov et al., 2023; Azar et al., 2024; Zhao et al., 2023; Meng et al., 2024; Wu et al., 2024b, a) have tried to bypass the need for an explicit reward model and RL learning, to directly optimize the policy model from preference data. Among them, DPO (Rafailov et al., 2023) is a milestone work that derives the optimal policy and substitutes it into the Bradley–Terry (BT) model (Bradley & Terry, 1952), which yields its objective:
@@ 3.1 Influence Function
where $v(\cdot)$ denotes a measurement to evaluate the model $\pi_{\theta}$, and $\theta_{\mathcal{D}}$ denotes the model trained on the dataset $\mathcal{D}$. Intuitively, computing the exact LOO effect requires training $|\mathcal{D}_{\text{train}}|+1$ separate models, which is infeasible in practice. To avoid repetitive retraining the model, influence function (IF) (Koh & Liang, 2017) provides a first-order Taylor approximation of the LOO effect using a gradient-based metric at the current parameters:
@@
where $\ell(\theta;d)$ is the training loss for the data $d$, and $H_{\theta}:=\nabla^{2}_{\theta}\mathcal{L}(\theta;\mathcal{D}_{\text{train}})$ is the Hessian matrix (Hampel, 1974) of the total training loss with respect to model parameters $\theta$. In practice, computing and inverting the Hessian matrix is often computationally intractable, especially for a large-scale model. A common approach is to approximate the Hessian $H_{\theta}$ as the identity matrix by assuming that the loss landscape is locally isotropic near $\theta$ (Koh & Liang, 2017; Wang et al., 2025; Xia et al., 2024). This simplification reduces the IF to a dot product between the gradient of the training data $d$ and the expected gradient of the validation set. Intuitively, a higher IF means the training gradient better aligns with the validation gradient, suggesting the data is more beneficial for generalization.
@@
where $g_{*}=\nabla_{\theta}\log\pi_{\theta}(y_{*}|x)$ denotes the gradient of the log-likelihood for response $y_{*}\in\{y_{w},y_{l}\}$, and $\Delta_{\theta}=\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}$ is the reward difference term in the DPO loss. Unlike IF under pointwise objectives, which assesses the gradient consistency of a give data point with validation set (Koh & Liang, 2017; Wang et al., 2025), the DPO-based IF focuses on the gradient difference consistency, i.e., $g_{w}-g_{l}$, with the validation preferences, thereby serving as a proxy to assess quality of a preference pair $d$. Vanilla IF implicitly assumes the validation set is an oracle of generalization, thereby equating high-IF data with high quality. However, under this assumption, IF may overfit to the specific validation set. The instantiation to SLiC loss is provided in Appendix B.1.
@@
3.2 Analysis: Truncated Influence Function (TIF)
Analytical Setup. We randomly sample a probing set $\mathcal{D}^{\text{prob}}=\{\mathcal{D}_{\text{train}}^{\text{prob}},\mathcal{D}_{\text{val}}^{\text{prob}},\mathcal{D}_{\text{test}}^{\text{prob}}\}$ from UltraFeedback (Cui et al., 2023), containing 5,000, 3,000, 300 preference pairs for training, validation, and testing, respectively. Qwen3-0.6B-Base (Yang et al., 2025) and Llama-3.2-1B (Vavekanand & Sam, 2024) are adopted as backbone models. Both models are first trained with supervised fine-tuning (SFT) on UltraChat-200K (Ding et al., 2023), followed by five epochs of DPO on $\mathcal{D}_{\text{train}}^{\text{prob}}$. At each-epoch DPO checkpoint, we compute the IF values of all training pairs using Eq. (5) with respect to $\mathcal{D}_{\text{val}}^{\text{prob}}$ and partition them into three equally sized subsets (small-, medium-, and large-IF). Each subset is then used to continue DPO training from the checkpoint, while we monitor the evaluation DPO loss and reward margin on $\mathcal{D}_{\text{test}}^{\text{prob}}$ to examine which type of data is more beneficial for alignment training.
4 Methodology: LossDiff-IRM Data Selection
@@ 4.1 Approximation Proxy of TIF
where $\pi_{\theta_{\text{val}}}$ is the auxiliary model aligned on validation set. An intuitive understanding is that a larger value indicates that moving from $\theta$ toward $\theta_{\text{val}}$ reduces the loss on $d$, which is consistent with the direction favored by the validation objective. We formally demonstrate that the LossDiff has a positive correlation with the IF in Appendix B.2. Empirically, the top panel of Figure 2 shows strong Pearson (Cohen et al., 2009) and Spearman (Hauke & Kossowski, 2011) correlations (e.g., $r=0.77$, $\rho=0.80$ on Qwen-0.6B-Base). Notably, this proxy requires only two forward passes per pair without backpropagation and remains tied to the validation set by construction.
@@ 5.1 Experimental Setup
Setup. Our experiments are conducted on diverse LLM families, including Llama-3.1-8B (Vavekanand & Sam, 2024), Qwen3-8B-Base (Yang et al., 2025), and the Pythia series (2.8B/1.4B/410M) (Biderman et al., 2023). Following prior work (Ko et al., 2024; Wu et al., 2024a), each pretrained LLM is first initialized with one epoch of supervised fine-tuning (SFT) on UltraChat-200k (Tunstall et al., 2024), which serves as the starting point for subsequent alignment. We adopt UltraFeedback-Binarized (Cui et al., 2023) as the alignment dataset, consistent with previous studies (Pattnaik et al., 2024; Gao et al., 2025; Ko et al., 2024). For validation set, we perform stratified sampling to select 20% of the alignment data according to the GPT-4 score margin between the chosen and rejected responses, and use the remaining 80% as training set. We employ DPO (Rafailov et al., 2023) and SLiC (Zhao et al., 2023) as core preference optimization algorithms used in our experiments. More details of experimental settings are provided in Appendix D.
@@
Evaluation Setup. We evaluate the aligned LLMs on open-ended generation tasks using both in-distribution (ID) and out-of-distribution (OOD) benchmarks. The ID evaluation is conducted on the UltraFeedback test split, while OOD evaluation includes AlpacaEval (Li et al., 2023) and Vicuna-Bench (Chiang et al., 2023). We compare against full-data training (Full Data) and preference data selection strategies commonly used in prior work (Pattnaik et al., 2024; Deng et al., 2025; Morimura et al., 2024): random sampling (Random), GPT-4 score filtering (GPT4), and external reward-model filtering (Reward Model). All methods are trained under the same setup for a fair comparison. We report two evaluation metrics via LLM-as-Judge (Zheng et al., 2023a): Single Score (Single) and Length-controlled Win Rate vs. SFT (WinRate). Further details are in Appendix D.3.
Methods
Dataset
Ratio
UltraFeedback AlpacaEval Vicuna-Bench Methods
Dataset
Ratio
@@ Methods
UltraFeedback AlpacaEval Vicuna-Bench
Llama-3.1-8B (DPO) Qwen3-8B-Base (DPO)
SFT 3.60 - 3.53 - 3.98 - SFT 6.97 - 6.88 - 7.94 -
Full Data 100% 5.77 77.61 5.87 78.41 6.04 73.75 Full Data 100% 7.64 61.41 7.92 63.85 8.21 62.14
Random 64% 5.52 74.83 5.59 75.93 5.46 68.13 Random 64% 7.71 61.47 7.94 64.12 8.26 58.93
GPT4 64% 6.04 80.57 6.21 81.09 6.86 80.31 GPT4 64% 7.69 62.19 8.01 63.81 8.28 52.19
Reward Model 64% 6.24 82.68 6.38 83.76 6.45 76.88 Reward Model 64% 7.81 64.19 8.24 69.35 8.56 66.25
LossDiff-IRM 64% 6.54 83.97 6.84 87.08 7.06 86.88 LossDiff-IRM 64% 8.05 67.32 8.36 71.52 8.72 67.19
Pythia-2.8B (DPO) Pythia-1.4B (DPO)
SFT 3.94 - 4.35 - 4.66 - SFT 3.50 - 3.65 - 4.20 -
Full Data 100% 4.60 70.53 4.95 67.05 5.35 67.50 Full Data 100% 3.70 65.88 3.99 66.17 4.71 64.38
Random 64% 4.54 68.27 4.79 64.03 5.15 68.13 Random 52% 3.78 67.43 4.05 68.16 4.56 61.25
GPT4 64% 4.71 72.02 4.96 67.10 5.39 73.75 GPT4 52% 3.96 70.75 4.28 70.96 4.89 68.75
Reward Model 64% 4.68 75.73 5.09 70.91 5.60 75.95 Reward Model 52% 3.83 70.80 4.18 70.83 4.84 68.75
LossDiff-IRM 64% 4.90 79.62 5.30 76.03 5.74 82.50 LossDiff-IRM 52% 4.23 78.49 4.49 76.43 5.28 76.88
Pythia-410M (DPO) Llama-3.1-8B (SLiC)
SFT 2.56 - 2.47 - 3.15 - SFT 3.60 - 3.53 - 3.98 -
Full Data 100% 2.81 75.25 2.77 73.51 3.10 69.37 Full Data 100% 5.09 70.72 5.13 72.13 5.40 71.88
Random 56% 2.94 76.03 2.92 76.62 3.58 70.63 Random 64% 4.94 69.52 4.89 67.05 5.26 67.50
GPT4 56% 2.95 76.78 2.95 79.81 3.49 79.37 GPT4 64% 5.48 75.61 5.40 72.89 6.05 67.81
Reward Model 56% 2.96 81.48 3.02 80.64 3.67 74.38 Reward Model 64% 5.34 73.64 5.54 75.03 5.55 68.75
LossDiff-IRM 56% 3.30 86.14 3.30 85.16 3.80 85.62 LossDiff-IRM 64% 5.94 79.51 5.84 78.84 5.85 76.56
Qwen3-8B-Base (SLiC) Pythia-2.8B (SLiC)
SFT 6.97 - 6.88 - 7.94 - SFT 3.94 - 4.35 - 4.66 -
Full Data 100% 7.55 59.54 7.61 59.71 8.05 54.69 Full Data 100% 4.36 67.46 4.48 61.66 4.90 65.00
Random 64% 7.57 57.76 7.63 62.05 7.97 47.94 Random 64% 4.31 63.00 4.56 59.73 5.16 65.62
GPT4 64% 7.64 59.91 7.89 63.62 8.21 58.37 GPT4 64% 4.50 69.09 4.73 63.00 5.12 65.62
Reward Model 64% 7.74 62.47 8.09 66.57 8.50 63.44 Reward Model 64% 4.43 70.03 4.76 64.92 5.50 71.88
LossDiff-IRM 64% 7.87 64.40 8.11 67.58 8.44 61.12 LossDiff-IRM 64% 4.82 76.36 5.02 68.85 5.47 74.38
Pythia-1.4B (SLiC) Pythia-410M (SLiC)
SFT 3.50 - 3.65 - 4.20 - SFT 2.56 - 2.47 - 3.15 -
Full Data 100% 3.66 63.58 3.98 63.68 4.67 60.00 Full Data 100% 2.81 73.80 2.82 73.91 3.31 71.25
Random 52% 3.81 66.18 3.96 63.99 4.34 56.25 Random 56% 2.67 69.23 2.69 70.68 3.09 63.12
GPT4 52% 3.84 69.24 4.12 65.55 4.69 63.75 GPT4 56% 2.80 72.75 2.85 74.53 3.23 75.62
Reward Model 52% 3.82 66.57 4.04 69.20 4.67 65.62 Reward Model 56% 2.87 77.41 2.94 78.39 3.50 75.00
LossDiff-IRM 52% 4.14 74.25 4.39 72.69 4.88 71.25 LossDiff-IRM 56% 3.07 80.08 3.09 84.39 3.83 79.36
Figure 3: Ablation study on WinRate. Comparison of LossDiff-IRM vs. its ablations that select data using only LossDiff (“w/ LossDiff”) or only IRM (“w/ IRM”). Top: DPO training. Bottom: SLiC training. Single-Score results are provided in Appendix E.1.

### v2

@@ Abstract
Large language model (LLM) alignment is typically achieved through learning from human preference comparisons, making the quality of preference data critical to its success. Existing studies often pre-process raw training datasets to identify valuable preference pairs using external reward models or off-the-shelf LLMs, achieving improved overall performance but rarely examining whether individual, selected data point is genuinely beneficial. We assess data quality through individual influence on validation data using our newly proposed truncated influence function (TIF), which mitigates the over-scoring present in traditional measures and reveals that preference data quality is inherently a property of the model. In other words, a data pair that benefits one model may harm another. This leaves the need to improve the preference data selection approaches to be adapting to specific models. To this end, we introduce two candidate scoring functions (SFs) that are computationally simpler than TIF and positively correlated with it. They are also model dependent and can serve as potential indicators of individual data quality for preference data selection. Furthermore, we observe that these SFs inherently exhibit errors when compared to TIF. To this end, we combine them to offset their diverse error sources, resulting in a simple yet effective data selection rule that enables the models to achieve a more precise selection of valuable preference data. We conduct experiments across diverse alignment benchmarks and various LLM families, with results demonstrating that better alignment performance can be achieved using less data, showing the generality of our findings and new methods. Our code is publicly available at https://github.com/tmlr-group/TIF_LossDiff-IRM.
@@ 1 Introduction
Reinforcement learning with human feedback (RLHF) has emerged as a dominant fine-tuning paradigm for aligning large language models (LLMs) with human preferences (Bai et al., 2022b; Wang et al., 2025b). Whether through training explicit reward models (Lambert et al., 2024) or optimizing policy with implicit rewards (Rafailov et al., 2023), the success of RLHF hinges heavily on the availability of high-quality preference data. Previous works (Shen et al., 2024; Pattnaik et al., 2024; Morimura et al., 2024; Deng et al., 2025; Chen et al., 2024) typically leverage external reward models or off-the-shelf LLMs to filter raw data, treating selected data as reliable sources for training their models or releasing them for open-source use (Cui et al., 2023; Bai et al., 2022a). This pre-processing paradigm dominates the RLHF community to improve data quality, contributing remarkably to the success of many publicly available LLMs such as Llama (Vavekanand and Sam, 2024), Qwen (Yang et al., 2025) and DeepSeek (Guo et al., 2025).
However, such pre-processing implicitly assumes that quality is an intrinsic property of data themselves: regardless of training configurations or models, certain data are consistently presumed to be more valuable for alignment than others. Therefore, a seemingly more realistic perspective is that the data quality is also a property of the model (Grosse et al., 2023; Xia et al., 2024). That is to say, some data points may be beneficial for certain models or configurations while being detrimental at others. We validate this new assumption as more reasonable based on the influence function (IF) (Koh and Liang, 2017), which quantifies the impact of each training data point on validation performance, thereby reflecting data quality. Nevertheless, we observe that the influence scores may overfit to the validation data, an issue that is particularly severe for open-world models like LLMs. To address this, we propose and verify a simple modification to the original IF, namely truncated influence function (TIF), which is shown to be more reliable for preference data selection and results in better overall performance. Using TIF, we verify that data quality is model-dependent, varying across different models, with some data points proving beneficial for certain models yet harmful at others.
@@ 2 Preliminary
Direct Preference Alignment. Traditional RLHF (Bai et al., 2022b) often follows a two-stage pipeline: first training a reward model on preference data, and then optimizing the policy LLM using reinforcement learning (RL) algorithm such as PPO (Schulman et al., 2017). The two-stage RLHF is relatively complicated and resource-intensive. Recently, many studies (Rafailov et al., 2023; Azar et al., 2024; Zhao et al., 2023; Meng et al., 2024; Wu et al., 2024b; a) have tried to bypass the need for an explicit reward model and RL learning, to directly optimize the policy model from preference data. Among them, DPO (Rafailov et al., 2023) is a milestone work that derives the optimal policy and substitutes it into the Bradley–Terry (BT) model (Bradley and Terry, 1952), which yields its objective:
@@ 3.1 Influence Function
where $v(\cdot)$ denotes a measurement to evaluate the model $\pi_{\theta}$, and $\theta_{\mathcal{D}}$ denotes the model trained on the dataset $\mathcal{D}$. Intuitively, computing the exact LOO effect requires training $|\mathcal{D}_{\text{train}}|+1$ separate models, which is infeasible in practice. To avoid repetitive retraining the model, influence function (IF) (Koh and Liang, 2017) provides a first-order Taylor approximation of the LOO effect using a gradient-based metric at the current parameters:
@@
where $\ell(\theta;d)$ is the training loss for the data $d$, and $H_{\theta}:=\nabla^{2}_{\theta}\mathcal{L}(\theta;\mathcal{D}_{\text{train}})$ is the Hessian matrix (Hampel, 1974) of the total training loss with respect to model parameters $\theta$. In practice, computing and inverting the Hessian matrix is often computationally intractable, especially for a large-scale model. A common approach is to approximate the Hessian $H_{\theta}$ as the identity matrix by assuming that the loss landscape is locally isotropic near $\theta$ (Koh and Liang, 2017; Wang et al., 2025a; Xia et al., 2024). This simplification reduces the IF to a dot product between the gradient of the training data $d$ and the expected gradient of the validation set. Intuitively, a higher IF means the training gradient better aligns with the validation gradient, suggesting the data is more beneficial for generalization.
@@
where $g_{*}=\nabla_{\theta}\log\pi_{\theta}(y_{*}|x)$ denotes the gradient of the log-likelihood for response $y_{*}\in\{y_{w},y_{l}\}$, and $\Delta_{\theta}=\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}$ is the reward difference term in the DPO loss. Unlike IF under pointwise objectives, which assesses the gradient consistency of a give data point with validation set (Koh and Liang, 2017; Wang et al., 2025a), the DPO-based IF focuses on the gradient difference consistency, i.e., $g_{w}-g_{l}$, with the validation preferences, thereby serving as a proxy to assess quality of a preference pair $d$. Vanilla IF implicitly assumes the validation set is an oracle of generalization, thereby equating high-IF data with high quality. However, under this assumption, IF may overfit to the specific validation set. The instantiation to SLiC loss is provided in Appendix B.1.
3.2 Analysis: Truncated Influence Function (TIF)
Analytical Setup. We randomly sample a probing set $\mathcal{D}^{\text{prob}}=\{\mathcal{D}_{\text{train}}^{\text{prob}},\mathcal{D}_{\text{val}}^{\text{prob}},\mathcal{D}_{\text{test}}^{\text{prob}}\}$ from UltraFeedback (Cui et al., 2023), containing 5,000, 3,000, 300 preference pairs for training, validation, and testing, respectively. Qwen3-0.6B-Base (Yang et al., 2025) and Llama-3.2-1B (Vavekanand and Sam, 2024) are adopted as backbone models. Both models are first trained with supervised fine-tuning (SFT) on UltraChat-200K (Ding et al., 2023), followed by five epochs of DPO on $\mathcal{D}_{\text{train}}^{\text{prob}}$. At each-epoch DPO checkpoint, we compute the IF values of all training pairs using Eq. (5) with respect to $\mathcal{D}_{\text{val}}^{\text{prob}}$ and partition them into three equally sized subsets (small-, medium-, and large-IF). Each subset is then used to continue DPO training from the checkpoint, while we monitor the evaluation DPO loss and reward margin on $\mathcal{D}_{\text{test}}^{\text{prob}}$ to examine which type of data is more beneficial for alignment training.
@@ 4.1 Approximation Proxy of TIF
where $\pi_{\theta_{\text{val}}}$ is the auxiliary model aligned on validation set. An intuitive understanding is that a larger value indicates that moving from $\theta$ toward $\theta_{\text{val}}$ reduces the loss on $d$, which is consistent with the direction favored by the validation objective. We formally demonstrate that the LossDiff has a positive correlation with the IF in Appendix B.2. Empirically, the top panel of Figure 2 shows strong Pearson (Cohen et al., 2009) and Spearman (Hauke and Kossowski, 2011) correlations (e.g., $r=0.77$, $\rho=0.80$ on Qwen-0.6B-Base). Notably, this proxy requires only two forward passes per pair without backpropagation and remains tied to the validation set by construction.
@@ 4.2 LossDiff-IRM Preference Data Selection
Methods Dataset Ratio UltraFeedback AlpacaEval Vicuna-Bench Arena-Hard Methods Dataset Ratio UltraFeedback AlpacaEval Vicuna-Bench Arena-Hard
Single $\uparrow$ WinRate $\uparrow$ Single $\uparrow$ WinRate $\uparrow$ Single $\uparrow$ WinRate $\uparrow$ Single $\uparrow$ WinRate $\uparrow$ Single $\uparrow$ WinRate $\uparrow$ Single $\uparrow$ WinRate $\uparrow$ Single $\uparrow$ WinRate $\uparrow$ Single $\uparrow$ WinRate $\uparrow$
Llama-3.1-8B (DPO) Qwen3-8B-Base (DPO)
SFT 3.60 - 3.53 - 3.98 - 2.63 - SFT 6.97 - 6.88 - 7.94 - 6.64 -
Full Data 100% 5.77 77.61 5.87 78.41 6.04 73.75 4.68 81.39 Full Data 100% 7.64 61.41 7.92 63.85 8.21 62.14 7.58 59.61
Random 64% 5.52 74.83 5.59 75.93 5.46 68.13 4.64 81.27 Random 64% 7.71 61.47 7.94 64.12 8.26 58.93 7.57 62.07
GPT4 64% 6.04 80.57 6.21 81.09 6.86 80.31 4.96 84.30 GPT4 64% 7.69 62.19 8.01 63.81 8.28 52.19 7.62 61.53
Reward Model 64% 6.24 82.68 6.38 83.76 6.45 76.88 5.13 86.19 Reward Model 64% 7.81 64.19 8.24 69.35 8.56 66.25 7.61 64.78
LossDiff-IRM 64% 6.54 83.97 6.84 87.08 7.06 86.88 5.59 88.40 LossDiff-IRM 64% 8.05 67.32 8.36 71.52 8.72 67.19 7.83 68.63
Pythia-2.8B (DPO) Pythia-1.4B (DPO)
SFT 3.94 - 4.35 - 4.66 - 2.74 - SFT 3.50 - 3.65 - 4.20 - 2.37 -
Full Data 100% 4.60 70.53 4.95 67.05 5.35 67.50 2.97 60.71 Full Data 100% 3.70 65.88 3.99 66.17 4.71 64.38 2.65 60.24
Random 64% 4.54 68.27 4.79 64.03 5.15 68.13 3.00 63.25 Random 52% 3.78 67.43 4.05 68.16 4.56 61.25 2.60 61.64
GPT4 64% 4.71 72.02 4.96 67.10 5.39 73.75 3.08 64.15 GPT4 52% 3.96 70.75 4.28 70.96 4.89 68.75 2.83 64.20
Reward Model 64% 4.68 75.73 5.09 70.91 5.60 75.95 3.03 63.03 Reward Model 52% 3.83 70.80 4.18 70.83 4.84 68.75 2.71 64.76
LossDiff-IRM 64% 4.90 79.62 5.30 76.03 5.74 82.50 3.26 71.64 LossDiff-IRM 52% 4.23 78.49 4.49 76.43 5.28 76.88 2.96 71.72
Pythia-410M (DPO) Llama-3.1-8B (SLiC)
SFT 2.56 - 2.47 - 3.15 - 1.91 - SFT 3.60 - 3.53 - 3.98 - 2.63 -
Full Data 100% 2.81 75.25 2.77 73.51 3.10 69.37 2.06 59.47 Full Data 100% 5.09 70.72 5.13 72.13 5.40 71.88 3.98 73.75
Random 56% 2.94 76.03 2.92 76.62 3.58 70.63 2.06 57.08 Random 64% 4.94 69.52 4.89 67.05 5.26 67.50 3.95 70.35
GPT4 56% 2.95 76.78 2.95 79.81 3.49 79.37 2.15 61.44 GPT4 64% 5.48 75.61 5.40 72.89 6.05 67.81 4.28 75.27
Reward Model 56% 2.96 81.48 3.02 80.64 3.67 74.38 2.16 60.59 Reward Model 64% 5.34 73.64 5.54 75.03 5.55 68.75 4.50 78.32
LossDiff-IRM 56% 3.30 86.14 3.30 85.16 3.80 85.62 2.38 69.63 LossDiff-IRM 64% 5.94 79.51 5.84 78.84 5.85 76.56 4.65 83.12
Qwen3-8B-Base (SLiC) Pythia-2.8B (SLiC)
SFT 6.97 - 6.88 - 7.94 - 6.64 - SFT 3.94 - 4.35 - 4.66 - 2.74 -
Full Data 100% 7.55 59.54 7.61 59.71 8.05 54.69 7.21 59.61 Full Data 100% 4.36 67.46 4.48 61.66 4.90 65.00 2.93 59.04
Random 64% 7.57 57.76 7.63 62.05 7.97 47.94 7.25 59.18 Random 64% 4.31 63.00 4.56 59.73 5.16 65.62 2.98 57.58
GPT4 64% 7.64 59.91 7.89 63.62 8.21 58.37 7.33 59.47 GPT4 64% 4.50 69.09 4.73 63.00 5.12 65.62 2.89 58.38
Reward Model 64% 7.74 62.47 8.09 66.57 8.50 63.44 7.38 62.27 Reward Model 64% 4.43 70.03 4.76 64.92 5.50 71.88 2.91 59.18
LossDiff-IRM 64% 7.87 64.40 8.11 67.58 8.44 61.12 7.61 62.20 LossDiff-IRM 64% 4.82 76.36 5.02 68.85 5.47 74.38 3.19 64.83
Pythia-1.4B (SLiC) Pythia-410M (SLiC)
SFT 3.50 - 3.65 - 4.20 - 2.37 - SFT 2.56 - 2.47 - 3.15 - 1.91 -
Full Data 100% 3.66 63.58 3.98 63.68 4.67 60.00 2.65 58.95 Full Data 100% 2.81 73.80 2.82 73.91 3.31 71.25 2.03 57.49
Random 52% 3.81 66.18 3.96 63.99 4.34 56.25 2.66 60.26 Random 56% 2.67 69.23 2.69 70.68 3.09 63.12 2.14 59.77
GPT4 52% 3.84 69.24 4.12 65.55 4.69 63.75 2.68 59.80 GPT4 56% 2.80 72.75 2.85 74.53 3.23 75.62 2.15 60.07
Reward Model 52% 3.82 66.57 4.04 69.20 4.67 65.62 2.67 61.95 Reward Model 56% 2.87 77.41 2.94 78.39 3.50 75.00 2.15 60.11
LossDiff-IRM 52% 4.14 74.25 4.39 72.69 4.88 71.25 2.85 67.76 LossDiff-IRM 56% 3.07 80.08 3.09 84.39 3.83 79.36 2.21 62.40
Figure 3: Ablation study on WinRate. Comparison of LossDiff-IRM vs. its ablations that select data using only LossDiff (“w/ LossDiff”) or only IRM (“w/ IRM”). Top: DPO training. Bottom: SLiC training. Single-Score results are provided in Appendix E.1 and Figure 12.
@@ 5.1 Experimental Setup
Setup. Our experiments are conducted on diverse LLM families, including Llama-3.1-8B (Vavekanand and Sam, 2024), Qwen3-8B-Base (Yang et al., 2025), and the Pythia series (2.8B/1.4B/410M) (Biderman et al., 2023). Following prior work (Ko et al., 2024; Wu et al., 2024a), each pretrained LLM is first initialized with one epoch of supervised fine-tuning (SFT) on UltraChat-200k (Tunstall et al., 2024), which serves as the starting point for subsequent alignment. We adopt UltraFeedback-Binarized (Cui et al., 2023) as the alignment dataset, consistent with previous studies (Pattnaik et al., 2024; Gao et al., 2025; Ko et al., 2024). For validation set, we perform stratified sampling to select 20% of the alignment data according to the GPT-4 score margin between the chosen and rejected responses, and use the remaining 80% as training set. We employ DPO (Rafailov et al., 2023) and SLiC (Zhao et al., 2023) as core preference optimization algorithms used in our experiments. More details of experimental settings are provided in Appendix D.
@@
Evaluation Setup. We evaluate the aligned LLMs on open-ended generation tasks using both in-distribution (ID) and out-of-distribution (OOD) benchmarks. The ID evaluation is conducted on the UltraFeedback test split, while OOD evaluation includes AlpacaEval (Li et al., 2023), Vicuna-Bench (Chiang et al., 2023), and Arena-Hard (Li et al., 2024). We compare against full-data training (Full Data) and several preference data selection strategies commonly used in prior work (Pattnaik et al., 2024; Deng et al., 2025; Morimura et al., 2024): random sampling (Random), GPT-4 score filtering (GPT4), and external reward-model filtering (Reward Model). We also compare with several existing methods: curriculum learning-based CurriDPO (Pattnaik et al., 2024), margin-based selection $M_{\text{AP}}$ (Huang et al., 2025), and rejection sampling-based preference data generation method RS-DPO (Khaki et al., 2024). All methods are trained under the same setup for a fair comparison. We report two evaluation metrics via LLM-as-Judge (Zheng et al., 2023a): Single Score (Single) and Length-controlled Win Rate vs. SFT (WinRate). Further details are in Appendix D.3.
Table 4: Performance comparisons of LossDiff-IRM with existing methods, including CurriDPO (Pattnaik et al., 2024), $M_{\text{AP}}$ (Huang et al., 2025), and RS-DPO (Khaki et al., 2024).
@@
Training Data UltraFeedback AlpacaEval Vicuna-Bench Arena-Hard
Single $\uparrow$ WinRate $\uparrow$ Single $\uparrow$ WinRate $\uparrow$ Single $\uparrow$ WinRate $\uparrow$ Single $\uparrow$ WinRate $\uparrow$
Llama-3.1-8B (DPO)
CurriDPO-GPT4 5.47 74.23 5.53 75.01 5.84 74.06 4.55 79.77
CurriDPO-Reward Model 5.51 74.62 5.54 74.29 5.59 74.06 4.62 79.49
$M_{\text{AP}}$ 6.04 79.99 6.21 80.88 6.34 74.69 5.08 85.30
RS-DPO 5.70 75.98 6.39 84.84 7.04 82.81 4.69 82.05
LossDiff-IRM 6.54 83.97 6.84 87.08 7.06 86.88 5.59 88.40
Qwen3-8B-Base (DPO)
CurriDPO-GPT4 7.61 61.04 7.74 62.35 8.16 54.69 7.52 62.51
CurriDPO-Reward Model 7.60 59.62 7.84 61.96 8.20 61.58 7.55 63.97
$M_{\text{AP}}$ 7.95 67.84 8.31 71.11 8.62 66.87 7.72 66.55
RS-DPO 7.88 64.87 8.36 74.07 8.93 71.90 7.48 61.32
LossDiff-IRM 8.05 67.32 8.36 71.52 8.72 67.19 7.83 68.63